# CLoMo: Counterfactual Logical Modification with Large Language Models

### v3

@@ 4 SES: Self-Evaluation Scores
Alignment with Human Evaluation We randomly selected 50 samples from the test set and examined how well the self-evaluation score (SES) matches human evaluation. Specifically, we use GPT-4 to implement SES. We recruit experts in argumentation to evaluate the modified arguments generated by GPT-4 on the selected 50 samples, scoring $1$ for good answers and 00 for bad answers. The Cohen’s Kappa coefficient between the human annotators is $\kappa=0.6785$, indicating substantial consistency of human perspective. We then use the self-evaluation score to evaluate the same group of modified arguments again. Cohen’s kappa coefficient between human and self-evaluation score is ${\kappa}=0.4391$. This indicates that the self-evaluation score is consistent with human evaluation. Therefore, we can use the self-evaluation score for more efficient logical modification evaluation. Moreover, since we can always check the logical reasoning ability of the model and its alignment with human evaluation, we can use a different model as the backbone to calculate this score.
Figure 3: Per-relation performances of LLaMA and LLaMA 2 families. NA: Necessary Assumption. SA: Sufficient Assumption. S: Strengthen. W: Weaken.
@@ 5.1 Main Results
We first evaluate the large language models GPT-3.5-Turbo OpenAI (2022) and GPT-4 OpenAI (2023). We also recruit 2 domain experts to contribute to the human performance. Table 3 demonstrates the results.
In the results, we see that the overall performance of the human experts on CLoMo is 0.542, indicating that CLoMo is quite a challenging task. In terms of the performance of the 4 logical relations, strengthen and weaken are relatively easy, while necessary assumption and sufficient assumption are more challenging for humans. In addition, the performance of GPT-4 is slightly lower than that of humans, but it also demonstrates strong counterfactual logical reasoning ability. And GPT-3.5-Turbo performs inferior to both human and GPT-4. Overall, large language models show great potential for counterfactual logical reasoning.
@@ 5.2 Fine-Tuning with Counterfactual Data
We then fine-tune LLaMA Touvron et al. (2023a) and LLaMA 2 Touvron et al. (2023b). We randomly split the CLoMo dataset into 60%/20%/20% training/development/test data, and fine-tune the models with the CLoMo training data. The implementation details are explained in Appendix C. We also compare the results with inference-only settings. The evaluation results are demonstrated in Figure 3.
@@
Table 4: Performances of smaller models in three inference-only settings. CoT: The chain-of-thought setting. Few: The few-shot setting. Zero: The zero-shot setting. More details are in Appendix A. NA: Necessary Assumption. SA: Sufficient Assumption. S: Strengthen. W: Weaken.
Figure 4: A successful case of counterfactual modification (reasoning type: Sufficient Assumption) generated by GPT-4, which makes a logically consistent ${}^{\prime}$ and is accordant with human reasoning.
Figure 5: An inferior case of counterfactual modification (reasoning type: Weaken) generated by GPT-4. It modifies Statement 1 by replacing intellectual activities with physical activities, where the logical restriction is not satisfied.
@@ 5.3 Performances of Small Language Models
We further test various language models on smaller scales and the results are shown in Table 4. All the models in Table 4 directly perform inference without further fine-tuning thus examining their original counterfactual abilities. The detailed settings are demonstrated in Appendix A, and the brief introductions of the language models are listed in Appendix B. We have the following findings: (1) Generally Speaking, all models perform inferiorly. Some of the models, for example, Baichuan2-7B-Chat in the few-shot setting, hardly solve any of the counterfactual questions. Among the models, Qwen-14B-Chat performs the best in both the chain-of-thought setting and the zero-shot setting. But all models perform inferior to large language models or human (2) The models perform better counterfactual reasoning with step-by-step reasoning processes (the CoT setting) while seeing more demonstrations in the prompt (the few-shot setting) harms the performances. It is indicated that the counterfactual cases have obscure reasoning patterns that are challenging for the models to transfer to unseen cases. (3) The performances are correlated to the model scales, but the gaps are not necessarily significant. For instance, Flan-T5-XXL (11B) performs generally better than Flan-T5-Large (780M), and Vicuna-13B-v1.5 performs better than Vicuna-7B-v1.5. (4) For the four different relations, it is shown that Sufficient Assumption and Weaken are significantly more challenging than the other two reasoning types. The possible reason is that compared to the other two reasoning types, Sufficient Assumption and Weaken require more reasoning steps such as reversed thinking. Two cases are further shown in Section 5.4. To sum up, the CLoMo task is challenging to current language models. Therefore, further investigation is needed on the counterfactual reasoning abilities of language models.
@@ 7 Limitations
The major limitation of this work is that we do not include multiple solutions as reference sentences in this version of CLoMo. However, the proposed SES score leverages LLM to conduct a human-aligned evaluation, so it reduces the impact of reference in evaluating the model performances.
8 Ethical Considerations
@@
Acknowledgements
The authors thank Guangyan Sun and Cen Li for reviewing the data.
Appendix A Implementation Details of Three Inference-Only Settings
Tables 5 to 16 demonstrate examples of three inference-only samples with input and output. The three inference-only settings are:
@@ Appendix A Implementation Details of Three Inference-Only Settings
Input:
Argument: Statement1: Philosopher : both the consequences and the motives of human actions have bearing on the moral worth of those actions. Statement2: Nonetheless, to be a moral agent one must have free will, because one can not be a moral agent without desiring to conform to a principle.
Premise1: Desiring to conform to a principle requires free will.
Premise2: It is impossible to have desires without also being a moral agent.
Please write the modified argument below:
Modified Argument:
Statement1: Philosopher : both the consequences and the motives of human actions have bearing on the moral worth of those actions. Statement2: Nonetheless, one can not have desires without desiring to conform to a principle, because one can not be a moral agent without desiring to conform to a principle.
Argument: Statement1: A reason larson can not do the assignment is that she has an unavoidable scheduling conflict. Statement2: On the other hand, a reason franks can not do the assignment is that he does not quite have the assertiveness the task requires. Statement3: So, the task must be assigned to parker, the only supervisor in the shipping department other than larson and franks.
Premise1: The task cannot be assigned to anyone other than a supervisor in the shipping department.
Premise2: The task cannot be assigned to anyone who has any kind of scheduling conflict.
Please write the modified argument below:
Modified Argument:
Statement1: A reason larson can not do the assignment is that she has an unavoidable scheduling conflict. Statement2: On the other hand, a reason franks can not do the assignment is that he does not quite have the assertiveness the task requires. Statement3: So, the task must be assigned to parker, the only one who do not have scheduling conflict in the shipping department other than larson and franks.
Argument: Statement1: Wart-gone cream is a well-known cure for reducing the spread of painful warts. A dime size amount rubbed on the infected area daily is the prescribed dosage, and applying the cream more often does not enhance the cream’s healing effects. Statement3: Hence, distributing coupons giving \$ 3.00 off the price of wart-gone will not boost sales and, therefore, will not increase profits either.
Premise1: The \$3. 00 coupons, if distributed, would not induce new customers, who might not have purchased the product otherwise, to try wart-gone.
Premise2: Wart-gone can be produced in small or large batches without appreciably lowering the pre-tube production costs.
Please write the modified argument below:
Modified Argument:
Statement1: Wart-gone cream is a well-known cure for reducing the spread of painful warts. A dime size amount rubbed on the infected area daily is the prescribed dosage, and applying the cream more often can enhance the cream’s healing effects. Statement3: Hence, distributing coupons giving \$ 3.00 off the price of wart-gone will boost produce and, however, will not change pre-tube production costs much.
Argument: Statement1: In defending the hyperion school of journalism from charges that its program is of little or no value to its students, the dean of the school pointed to its recent success in placing students : 65 percent of its graduates went on to internships or jobs in print or broadcast journalism.
Premise1: More than half of the school’s students came from jobs in journalism to improve their skills.
Premise2: Some newspaper editors do not regard journalism schools as a necessary part of the training of a journalist.
Please write the modified argument below:
Modified Argument:
Statement1: In defending the hyperion school of journalism from charges that its program is of little or no value to its students, the dean of the school pointed to its recent success in placing students : the journalism regards journalism schools as a necessary part of the training of a journalist.
Argument: Statement1: Consumer advocate : there is no doubt that the government is responsible for the increased cost of gasoline, because the government’s policies have significantly increased consumer demand for fuel, and as a result of increasing demand, the price of gasoline has risen steadily.
Premise1: The government can bear responsibility for that which it indirectly causes.
Premise2: Consumer demand for gasoline cannot increase without causing gasoline prices to increase.
Please write the modified argument below:
Output:
Input:
Argument: Statement1: Philosopher : both the consequences and the motives of human actions have bearing on the moral worth of those actions. Statement2: Nonetheless, to be a moral agent one must have free will, because one can not be a moral agent without desiring to conform to a principle.
Premise1: Desiring to conform to a principle requires free will.
Premise2: It is impossible to have desires without also being a moral agent.
Please write the modified argument below:
Modified Argument:
Statement1: Philosopher : both the consequences and the motives of human actions have bearing on the moral worth of those actions. Statement2: Nonetheless, one can not have desires without desiring to conform to a principle, because one can not be a moral agent without desiring to conform to a principle.
Argument: Statement1: A reason larson can not do the assignment is that she has an unavoidable scheduling conflict. Statement2: On the other hand, a reason franks can not do the assignment is that he does not quite have the assertiveness the task requires. Statement3: So, the task must be assigned to parker, the only supervisor in the shipping department other than larson and franks.
Premise1: The task cannot be assigned to anyone other than a supervisor in the shipping department.
Premise2: The task cannot be assigned to anyone who has any kind of scheduling conflict.
Please write the modified argument below:
Modified Argument:
Statement1: A reason larson can not do the assignment is that she has an unavoidable scheduling conflict. Statement2: On the other hand, a reason franks can not do the assignment is that he does not quite have the assertiveness the task requires. Statement3: So, the task must be assigned to parker, the only one who do not have scheduling conflict in the shipping department other than larson and franks.
Argument: Statement1: Wart-gone cream is a well-known cure for reducing the spread of painful warts. A dime size amount rubbed on the infected area daily is the prescribed dosage, and applying the cream more often does not enhance the cream’s healing effects. Statement3: Hence, distributing coupons giving \$ 3.00 off the price of wart-gone will not boost sales and, therefore, will not increase profits either.
Premise1: The \$3. 00 coupons, if distributed, would not induce new customers, who might not have purchased the product otherwise, to try wart-gone.
Premise2: Wart-gone can be produced in small or large batches without appreciably lowering the pre-tube production costs.
Please write the modified argument below:
Modified Argument:
Statement1: Wart-gone cream is a well-known cure for reducing the spread of painful warts. A dime size amount rubbed on the infected area daily is the prescribed dosage, and applying the cream more often can enhance the cream’s healing effects. Statement3: Hence, distributing coupons giving \$ 3.00 off the price of wart-gone will boost produce and, however, will not change pre-tube production costs much.
Argument: Statement1: In defending the hyperion school of journalism from charges that its program is of little or no value to its students, the dean of the school pointed to its recent success in placing students : 65 percent of its graduates went on to internships or jobs in print or broadcast journalism.
Premise1: More than half of the school’s students came from jobs in journalism to improve their skills.
Premise2: Some newspaper editors do not regard journalism schools as a necessary part of the training of a journalist.
Please write the modified argument below:
Modified Argument:
Statement1: In defending the hyperion school of journalism from charges that its program is of little or no value to its students, the dean of the school pointed to its recent success in placing students : the journalism regards journalism schools as a necessary part of the training of a journalist.
Argument: Statement1: Super express shipping company has implemented a new distribution system that can get almost every package to its destination the day after it is sent. The company worries that this more efficient system will result in lower sales of its premium next-day delivery service, because its two day service will usually arrive the following day anyway. Statement2: The company plans to encourage sales of its next-day service by intentionally delaying delivery of its two-day packages so that they will not be delivered the following day, even if the package arrives at its destination city in time for next-day delivery.
Premise1: Deliberate delay of packages will not affect the company’s image in a way that significantly reduces its ability to attract and retain customers.
Premise2: The overnight service is too expensive to be attractive to most customers currently.
Please write the modified argument below:
Output:
Input:
Argument: Statement1: Philosopher : both the consequences and the motives of human actions have bearing on the moral worth of those actions. Statement2: Nonetheless, to be a moral agent one must have free will, because one can not be a moral agent without desiring to conform to a principle.
Premise1: Desiring to conform to a principle requires free will.
Premise2: It is impossible to have desires without also being a moral agent.
Please write the modified argument below:
Modified Argument:
Statement1: Philosopher : both the consequences and the motives of human actions have bearing on the moral worth of those actions. Statement2: Nonetheless, one can not have desires without desiring to conform to a principle, because one can not be a moral agent without desiring to conform to a principle.
Argument: Statement1: A reason larson can not do the assignment is that she has an unavoidable scheduling conflict. Statement2: On the other hand, a reason franks can not do the assignment is that he does not quite have the assertiveness the task requires. Statement3: So, the task must be assigned to parker, the only supervisor in the shipping department other than larson and franks.
Premise1: The task cannot be assigned to anyone other than a supervisor in the shipping department.
Premise2: The task cannot be assigned to anyone who has any kind of scheduling conflict.
Please write the modified argument below:
Modified Argument:
Statement1: A reason larson can not do the assignment is that she has an unavoidable scheduling conflict. Statement2: On the other hand, a reason franks can not do the assignment is that he does not quite have the assertiveness the task requires. Statement3: So, the task must be assigned to parker, the only one who do not have scheduling conflict in the shipping department other than larson and franks.
Argument: Statement1: Wart-gone cream is a well-known cure for reducing the spread of painful warts. A dime size amount rubbed on the infected area daily is the prescribed dosage, and applying the cream more often does not enhance the cream’s healing effects. Statement3: Hence, distributing coupons giving \$ 3.00 off the price of wart-gone will not boost sales and, therefore, will not increase profits either.
Premise1: The \$3. 00 coupons, if distributed, would not induce new customers, who might not have purchased the product otherwise, to try wart-gone.
Premise2: Wart-gone can be produced in small or large batches without appreciably lowering the pre-tube production costs.
Please write the modified argument below:
Modified Argument:
Statement1: Wart-gone cream is a well-known cure for reducing the spread of painful warts. A dime size amount rubbed on the infected area daily is the prescribed dosage, and applying the cream more often can enhance the cream’s healing effects. Statement3: Hence, distributing coupons giving \$ 3.00 off the price of wart-gone will boost produce and, however, will not change pre-tube production costs much.
Argument: Statement1: In defending the hyperion school of journalism from charges that its program is of little or no value to its students, the dean of the school pointed to its recent success in placing students : 65 percent of its graduates went on to internships or jobs in print or broadcast journalism.
Premise1: More than half of the school’s students came from jobs in journalism to improve their skills.
Premise2: Some newspaper editors do not regard journalism schools as a necessary part of the training of a journalist.
Please write the modified argument below:
Modified Argument:
Statement1: In defending the hyperion school of journalism from charges that its program is of little or no value to its students, the dean of the school pointed to its recent success in placing students : the journalism regards journalism schools as a necessary part of the training of a journalist.
Argument: Statement1: Often patients with ankle fractures that are stable, and thus do not require surgery, are given follow-up x-rays because their orthopedists are concerned about possibly having misjudged the stability of the fracture. Statement2: When a number of follow-up x-rays were reviewed, however, all the fractures that had initially been judged stable were found to have healed correctly. Statement3: Therefore, it is a waste of money to order follow-up x-rays of ankle fracture initially judged stable.
Premise1: X-rays of patients of many different orthopedists working in several hospitals were reviewed.
Premise2: Doctors who are general practitioners rather than orthopedists are less likely than orthopedists to judge the stability of an ankle fracture correctly.
Please write the modified argument below:
Output:
Input:
Argument: Statement1: Philosopher : both the consequences and the motives of human actions have bearing on the moral worth of those actions. Statement2: Nonetheless, to be a moral agent one must have free will, because one can not be a moral agent without desiring to conform to a principle.
Premise1: Desiring to conform to a principle requires free will.
Premise2: It is impossible to have desires without also being a moral agent.
Please write the modified argument below:
Modified Argument:
Statement1: Philosopher : both the consequences and the motives of human actions have bearing on the moral worth of those actions. Statement2: Nonetheless, one can not have desires without desiring to conform to a principle, because one can not be a moral agent without desiring to conform to a principle.
Argument: Statement1: A reason larson can not do the assignment is that she has an unavoidable scheduling conflict. Statement2: On the other hand, a reason franks can not do the assignment is that he does not quite have the assertiveness the task requires. Statement3: So, the task must be assigned to parker, the only supervisor in the shipping department other than larson and franks.
Premise1: The task cannot be assigned to anyone other than a supervisor in the shipping department.
Premise2: The task cannot be assigned to anyone who has any kind of scheduling conflict.
Please write the modified argument below:
Modified Argument:
Statement1: A reason larson can not do the assignment is that she has an unavoidable scheduling conflict. Statement2: On the other hand, a reason franks can not do the assignment is that he does not quite have the assertiveness the task requires. Statement3: So, the task must be assigned to parker, the only one who do not have scheduling conflict in the shipping department other than larson and franks.
Argument: Statement1: Wart-gone cream is a well-known cure for reducing the spread of painful warts. A dime size amount rubbed on the infected area daily is the prescribed dosage, and applying the cream more often does not enhance the cream’s healing effects. Statement3: Hence, distributing coupons giving \$ 3.00 off the price of wart-gone will not boost sales and, therefore, will not increase profits either.
Premise1: The \$3. 00 coupons, if distributed, would not induce new customers, who might not have purchased the product otherwise, to try wart-gone.
Premise2: Wart-gone can be produced in small or large batches without appreciably lowering the pre-tube production costs.
Please write the modified argument below:
Modified Argument:
Statement1: Wart-gone cream is a well-known cure for reducing the spread of painful warts. A dime size amount rubbed on the infected area daily is the prescribed dosage, and applying the cream more often can enhance the cream’s healing effects. Statement3: Hence, distributing coupons giving \$ 3.00 off the price of wart-gone will boost produce and, however, will not change pre-tube production costs much.
Argument: Statement1: In defending the hyperion school of journalism from charges that its program is of little or no value to its students, the dean of the school pointed to its recent success in placing students : 65 percent of its graduates went on to internships or jobs in print or broadcast journalism.
Premise1: More than half of the school’s students came from jobs in journalism to improve their skills.
Premise2: Some newspaper editors do not regard journalism schools as a necessary part of the training of a journalist.
Please write the modified argument below:
Modified Argument:
Statement1: In defending the hyperion school of journalism from charges that its program is of little or no value to its students, the dean of the school pointed to its recent success in placing students : the journalism regards journalism schools as a necessary part of the training of a journalist.
Argument: Statement1: To suit the needs of corporate clients, advertising agencies have successfully modified a strategy originally developed for political campaigns. Statement2: This strategy aims to provide clients with free publicity and air time by designing an advertising campaign that is controversial, thus drawing prime-time media coverage and evoking public comment by officials.
Premise1: The usefulness of an advertising campaign is based solely on the degree to which the campaign’s advertisements persuade their audiences.
Premise2: Advertising agencies that specialize in campaigns for corporate clients are not usually chosen for political campaigns.
Please write the modified argument below:
Output:
Input:
Argument: Statement1: Consumer advocate : there is no doubt that the government is responsible for the increased cost of gasoline, because the government’s policies have significantly increased consumer demand for fuel, and as a result of increasing demand, the price of gasoline has risen steadily.
Premise1: The government can bear responsibility for that which it indirectly causes.
Premise2: Consumer demand for gasoline cannot increase without causing gasoline prices to increase.
Please write the modified argument below:
Output:
Input:
Argument: Statement1: Super express shipping company has implemented a new distribution system that can get almost every package to its destination the day after it is sent. The company worries that this more efficient system will result in lower sales of its premium next-day delivery service, because its two day service will usually arrive the following day anyway. Statement2: The company plans to encourage sales of its next-day service by intentionally delaying delivery of its two-day packages so that they will not be delivered the following day, even if the package arrives at its destination city in time for next-day delivery.
Premise1: Deliberate delay of packages will not affect the company’s image in a way that significantly reduces its ability to attract and retain customers.
Premise2: The overnight service is too expensive to be attractive to most customers currently.
Please write the modified argument below:
Output:
Input:
Argument: Statement1: Often patients with ankle fractures that are stable, and thus do not require surgery, are given follow-up x-rays because their orthopedists are concerned about possibly having misjudged the stability of the fracture. Statement2: When a number of follow-up x-rays were reviewed, however, all the fractures that had initially been judged stable were found to have healed correctly. Statement3: Therefore, it is a waste of money to order follow-up x-rays of ankle fracture initially judged stable.
Premise1: X-rays of patients of many different orthopedists working in several hospitals were reviewed.
Premise2: Doctors who are general practitioners rather than orthopedists are less likely than orthopedists to judge the stability of an ankle fracture correctly.
Please write the modified argument below:
Output:
Input:
Argument: Statement1: To suit the needs of corporate clients, advertising agencies have successfully modified a strategy originally developed for political campaigns. Statement2: This strategy aims to provide clients with free publicity and air time by designing an advertising campaign that is controversial, thus drawing prime-time media coverage and evoking public comment by officials.
Premise1: The usefulness of an advertising campaign is based solely on the degree to which the campaign’s advertisements persuade their audiences.
Premise2: Advertising agencies that specialize in campaigns for corporate clients are not usually chosen for political campaigns.
Please write the modified argument below:
Output:
Input:
Argument: Statement1: Consumer advocate : there is no doubt that the government is responsible for the increased cost of gasoline, because the government’s policies have significantly increased consumer demand for fuel, and as a result of increasing demand, the price of gasoline has risen steadily.
Premise1: The government can bear responsibility for that which it indirectly causes.
Premise2: Consumer demand for gasoline cannot increase without causing gasoline prices to increase.
Please write the modified argument below. Let’s think step by step:
Output:
Input:
Argument: Statement1: Super express shipping company has implemented a new distribution system that can get almost every package to its destination the day after it is sent. The company worries that this more efficient system will result in lower sales of its premium next-day delivery service, because its two day service will usually arrive the following day anyway. Statement2: The company plans to encourage sales of its next-day service by intentionally delaying delivery of its two-day packages so that they will not be delivered the following day, even if the package arrives at its destination city in time for next-day delivery.
Premise1: Deliberate delay of packages will not affect the company’s image in a way that significantly reduces its ability to attract and retain customers.
Premise2: The overnight service is too expensive to be attractive to most customers currently.
Please write the modified argument below. Let’s think step by step:
Output:
Input:
Argument: Statement1: Often patients with ankle fractures that are stable, and thus do not require surgery, are given follow-up x-rays because their orthopedists are concerned about possibly having misjudged the stability of the fracture. Statement2: When a number of follow-up x-rays were reviewed, however, all the fractures that had initially been judged stable were found to have healed correctly. Statement3: Therefore, it is a waste of money to order follow-up x-rays of ankle fracture initially judged stable.
Premise1: X-rays of patients of many different orthopedists working in several hospitals were reviewed.
Premise2: Doctors who are general practitioners rather than orthopedists are less likely than orthopedists to judge the stability of an ankle fracture correctly.
Please write the modified argument below. Let’s think step by step:
Output:
Input:
Argument: Statement1: To suit the needs of corporate clients, advertising agencies have successfully modified a strategy originally developed for political campaigns. Statement2: This strategy aims to provide clients with free publicity and air time by designing an advertising campaign that is controversial, thus drawing prime-time media coverage and evoking public comment by officials.
Premise1: The usefulness of an advertising campaign is based solely on the degree to which the campaign’s advertisements persuade their audiences.
Premise2: Advertising agencies that specialize in campaigns for corporate clients are not usually chosen for political campaigns.
Please write the modified argument below. Let’s think step by step:
Output:
Appendix B Compared Models

### v4

@@ 4 SES: Self-Evaluation Scores
[…]
Alignment with Human Evaluation We randomly select 50 samples from the test set and examined how well the self-evaluation score (SES) matches human evaluation. Specifically, we use GPT-4 to implement SES. We recruit experts in argumentation to evaluate the modified arguments generated by GPT-4 on the selected 50 samples, scoring $1$ for good answers and 00 for bad answers. The Cohen’s Kappa coefficient between the human annotators is $\kappa=0.6785$, indicating substantial consistency of human perspective. We then use the self-evaluation score to evaluate the same group of modified arguments again. Cohen’s kappa coefficient between human and self-evaluation score is ${\kappa}=0.4391$. This indicates that the self-evaluation score is a safe reference and assistance for humans as the first study on automating the evaluation of the highly challenging counterfactual modification task. Therefore, we can apply the SES score for more efficient logical modification evaluation. Also, we believe in a further improved automated score for this task, which we leave as a future work. Additionally, the SES score can adjust to the latest and most cutting-edge models as backbones and it is convenient to check backbones’ logical reasoning ability and alignment with humans.
@@ 5.1 Main Results
We first evaluate the large language models GPT-3.5-Turbo OpenAI (2022), GPT-4 OpenAI (2023), and the latest GPT-4o11footnotemark: 1. We also recruit 2 domain experts to contribute to the human performance. Table 5 demonstrates the results. The backbone of SES is GPT-4.
The overall performance of the human experts on CLoMo is 0.580, indicating that CLoMo is quite a challenging task. Among the 4 logical relations, necessary assumption and sufficient assumption are more challenging for humans. We consider a model to have counterfactual reasoning capabilities if its SES is comparable to, or even exceeds, those of humans. The performance of GPT-4 is slightly lower than that of humans, but it also demonstrates strong counterfactual logical reasoning ability. And GPT-3.5-Turbo performs inferior to both human and GPT-4. We notice that GPT-4o exceeds human performance except that GPT-4o still has difficulty reasoning with the Weaken relation. Overall, large language models show great potential for counterfactual logical reasoning.
Figure 3: Per-relation performances of fine-tuned and inference-only LLaMA and LLaMA 2 families. NA: Necessary Assumption. SA: Sufficient Assumption. S: Strengthen. W: Weaken.
CoT Few Zero
R w/o R Full w/o R Full w/o R Full
Necessary Assumption 0.434 0.494 0.052 0.633 0.234 0.557
Sufficient Assumption 0.286 0.500 0.000 0.357 0.234 0.557
Strengthen 0.457 0.771 0.114 0.800 0.343 0.771
Weaken 0.222 0.611 0.014 0.750 0.243 0.708
Table 6: LLaMA 2-7B performances on test data with unseen logical relation. w/o R: LLaMA 2-7B fine-tuned without R-type data. Full: LLaMA 2-7B fine-tuned with full training data.
@@ 5.2 Fine-Tuning with Counterfactual Data
We then fine-tune LLaMA Touvron et al. (2023a) and LLaMA 2 Touvron et al. (2023b). We randomly split the CLoMo dataset into 60%/20%/20% training/development/test data, and fine-tune the models with the CLoMo training data. The implementation details are explained in Appendix F. We also compare the results with inference-only settings. The evaluation results are demonstrated in Figure 3.
@@
Ablation Study To study if there are only some CLoMo training data is sufficient to reveal unseen logical relations, we further fine-tuned LLaMA-7b on training data excluding logical relation R (R=Necessary Assumption/Sufficient Assumption/Strengthen/Weaken), and evaluated it on test data that includes R. We then compared the performance of this model with itself trained on the full data set. Table 6 shows that the performance of the unseen logical relation does drop drastically. This suggests that comprehensive learning across all types of logical relations is crucial.
@@
Table 7: Performances of smaller models in three inference-only settings. CoT: The chain-of-thought setting. Few: The few-shot setting. Zero: The zero-shot setting. More details are in Appendix D. NA: Necessary Assumption. SA: Sufficient Assumption. S: Strengthen. W: Weaken.
Figure 4: A successful case of counterfactual modification (reasoning type: Sufficient Assumption) generated by GPT-4, which makes a logically consistent Argument and is accordant with human reasoning. Counterfactually modified segments are underlined.
Figure 5: An inferior case of counterfactual modification (reasoning type: Weaken) generated by GPT-4. It modifies Statement 1 by replacing intellectual activities with physical activities, where the logical restriction is not satisfied. Counterfactually modified segments are underlined.
@@ 5.3 Performances of Small Language Models
We further test various language models on smaller scales and the results are shown in Table 7. All the models in Table 7 directly perform inference without further fine-tuning thus examining their original counterfactual abilities. The detailed settings are demonstrated in Appendix D, and the brief introductions of the language models are listed in Appendix E. We have the following findings: (1) Generally Speaking, all models perform inferiorly. Some of the models, for example, Baichuan2-7B-Chat in the few-shot setting, hardly solve any of the counterfactual questions. Among the models, Qwen-14B-Chat performs the best in both the chain-of-thought setting and the zero-shot setting. But all models perform inferior to large language models or human (2) The models perform better counterfactual reasoning with step-by-step reasoning processes (the CoT setting) while seeing more demonstrations in the prompt (the few-shot setting) harms the performances. It is indicated that the counterfactual cases have obscure reasoning patterns that are challenging for the models to transfer to unseen cases. (3) The performances are correlated to the model scales, but the gaps are not necessarily significant. For instance, Flan-T5-XXL (11B) in general performs better than Flan-T5-Large (780M), and Vicuna-13B-v1.5 performs better than Vicuna-7B-v1.5. (4) For the four different relations, it is shown that Sufficient Assumption and Weaken are significantly more challenging than the other two reasoning types. The possible reason is that compared to the other two reasoning types, Sufficient Assumption and Weaken require more reasoning steps such as reversed thinking. Two cases are further shown in Section 5.4. To sum up, the CLoMo task is challenging to current language models. Therefore, further investigation is needed on the counterfactual reasoning abilities of language models.
@@ 7 Limitations
The major limitation of this work is that we do not include multiple reference sentences in this version of CLoMo. However, the proposed SES score leverages LLM to conduct a human-aligned evaluation, so it reduces the impact of reference in evaluating the model performances.
Acknowledgements
This work was supported in part by the National Science and Technology Major Project under Grant No. 2020AAA0109704, the Research Grants Council of the Hong Kong SAR under Grant GRF 11217823 and Collaborative Research Fund C1042-23GF, the National Natural Science Foundation of China under Grant 62371411, InnoHK initiative, the Government of the HKSAR, Laboratory for AI-Powered Financial Technologies. The authors thank Dr. Xingchi Su, Guangyan Sun, and Cen Li for their great effort in carefully reviewing the data.
Appendix A Ethical Considerations
@@
Appendix B Samples from CLoMo
Tables 8 and 10 demonstrate examples from the CLoMo dataset.
Appendix C Data Source
Since applicable data for the proposed task is lacking, we build a benchmark dataset by carefully selecting argument texts and collecting human annotation of the modified Argument. We choose to use ReClor Yu et al. (2020) as the source data, considering that ReClor includes standardized multiple-choice questions on argument texts and logical relations from LSAT. We then recruit domain experts to conduct the annotation.
Appendix D Implementation Details of Three Inference-Only Settings
Tables 11 to 22 demonstrate examples of three inference-only samples with input and output. The three inference-only settings are:
@@ Appendix D Implementation Details of Three Inference-Only Settings
Appendix E Compared Models